# Compact stellar systems hosting an intermediate mass black hole: magnetohydrodynamic study of inflow-outflow dynamics

### v1

@@ 2 Methods
We conduct our simulations with athena++, a 3D grid–based scheme that solves the equations of conservative MHD (Stone et al., 2020). Our simulations are based on the HD stellar wind model described in Ressler et al. (2018) and extended in Ressler et al. (2019a, b) to include magnetized winds. The stellar winds are treated as source terms in mass, momentum, energy, and magnetic field moving on fixed Keplerian orbits. The hydrodynamic properties of the stellar winds are characterized by their mass–loss rates, $\dot{M_{\mathrm{w}}}$, and terminal stellar-wind velocities, $v_{\mathrm{w}}$. Furthermore, we include magnetization of the winds in some of our simulations. The magnetic fields are solely toroidal with respect to the spin axes of the stars (unknown but chosen randomly), and their magnitude is expressed through the $\beta_{\mathrm{w}}$ parameter, defined as the ratio between the wind ram pressure and its magnetic pressure at the equator. In the general scenario involving magnetized winds, the simulation solves the following set of equations:
@@
$\frac{\partial\mathbf{B}}{\partial t}-\nabla\times(\mathbf{v}\times\mathbf{B})$ =×$=\nabla\times(\mathbf{\tilde{E}}_{\mathrm{w}}),$
[…]
@@ 2.1 Cooling function and chemical abundance models
Due to the close proximity of the stellar winds, they interact and shock-heat, resulting in significant radiative losses due to optically thin bremsstrahlung and line cooling. To calculate the cooling function, $\Lambda$, for arbitrary hydrogen, helium, and metal mass fractions ($X$, $Y$, $Z$ respectively), we first determine the cooling curve for the solar abundances $X_{\odot}=0.7491,\,Y_{\odot}=0.2377,\,Z_{\odot}=0.0133$, presented in Lodders (2003). This is done using the spectral analysis code spex (Kaastra et al., 1996), following the methodology outlined in Schure et al. (2009) to calculate the separate contributions of individual elements to the cooling function,
@@
Here, we assume that oxygen contributes the majority of the mass in metals and treat both $\mu_{\rm e}$ and $\mu$ as constants. The cooling function $\Lambda$ was implemented by calculating the cooling rate per unit volume using a piece-wise power law approximation over a range of temperatures from $10^{4}$ to $10^{9}$ K. For additional details on the computation and implementation of the cooling function in the simulations, refer to Sect. 2.2 in Ressler et al. (2018).
The real chemical composition of WR stars in compact stellar clusters remains largely uncertain. We, therefore, use three different chemical abundance models of WR stars used previously in Galactic center simulations (Cuadra et al., 2008; Calderón et al., 2016; Russell et al., 2017). First, we consider hydrogen-depleted stars ($X=0$) with higher metallicity ($Z=3Z_{\odot}$), referred to as Model I in Table 1. This model is motivated by the fact that WR stars typically lose their outer hydrogen later in previous stages of stellar evolution. Additionally, we explore two alternative chemical abundance models: one with a higher hydrogen content (Model II), which aligns with a realistic expectation that the stars are not completely starved of hydrogen, and another with significantly elevated metallicity (Model III), to investigate effects such as enhanced cooling efficiency and the potential formation of cool clumps due to stellar wind interactions. The accretion flow in Galactic center simulations can be quite sensitive to these mass fractions, as was recently shown by Calderón et al. (2025).
@@ 2.2 Simulation Setup
Our setup is loosely based on the stellar association IRS 13E within the Milky Way nuclear stellar cluster in the close vicinity of Sgr A*. The six WR stars in our simulation are placed at fixed distances from the central black hole, matching the observed projected distances of these stars from the hypothetical IMBH with an approximate mass of $\sim 3\times 10^{4}M_{\odot}$, associated with the IRS 13E3 source (see Peißker et al., 2024, for more details). That said, we use IRS 13E primarily as a reference for the stellar properties, taking a more flexible approach to other key properties, such as the spatial distribution of the stars, their chemical abundance models, and the inclusion of wind magnetization.
@@
Simulation Chemical abundance model Stellar distribution $\beta_{\mathrm{w}}$ Resolution SMR levels
1 I spherical A $\infty$ $128^{3}$ 5
2 II spherical B $\infty$ $256^{3}$ 5
3 II spherical B 100 $256^{3}$ 5
4 III spherical A $\infty$ $256^{3}$ 5
5 I disk-like $\infty$ $128^{3}$ 5
6 III disk-like $\infty$ $256^{3}$ 5
  • Notes. A and B in the spherical stellar distributions denote different random realizations of the system.
Table 3: Distances from the IMBH, orbital velocities, and orbital periods of the stars in our simulations. The setup is inspired by the IRS 13E stellar association in the Milky Way nuclear stellar cluster, and the listed distances are derived from the projected distances of each star from the putative IMBH associated with IRS 13E3. Source IDs are adopted from Peißker et al. (2023). a
Source ID $r$ [$10^{-3}$pc] $v$ [$\text{km\,s}^{-1}$] $P$ [yr]
@@ 6 III disk-like $\infty$ $256^{3}$ 5
[…]
Figure 1: 3D volume rendering of density of the simulated stellar cluster. Left: Colliding stellar winds 620 years into the Model II spherical simulation. Right: Outbursts of cool clumps 815 years into the Model III spherical simulation, driven by the gas’s higher metal content, which enhances cooling efficiency and amplifies the development of the thin-shell instability.
The setup summarized in Table 3 indicates that for the mean stellar distance of $r\sim 10\,{\rm mpc}$, the orbital velocity is $v=(GM_{\bullet}/r)^{1/2}\simeq 113.6(M_{\bullet}/3\times 10^{4}\,M_{\odot})^{1/2}(r/10\,{\rm mpc})^{-1/2}\,{\rm km\,s^{-1}}$. This implies a stellar-wind velocity to orbital velocity ratio of $v_{\rm w}/v\sim 8.8$. For the stars bound to Sgr A*, the orbital velocities within the clockwise disc are of the order of $v\sim 415\,(M_{\bullet}/4\times 10^{6}\,M_{\odot})^{1/2}(r/0.1\,{\rm pc})^{-1/2}\,{\rm km\,s^{-1}}$ and the ratio is $v_{\rm w}/v\sim 2.4$, which is $\sim 4$ times smaller than for the case of adopted orbits around the IMBH. The difference in $v_{\rm w}/v$ clearly motivates the exploration of the wind-fed accretion regime in compact stellar systems around the IMBH.
Another motivation is the much higher mean stellar density in comparison with young massive stars around Sgr A*. Around Sgr A*, there are about $\sim 200$ young massive stars within $r\lesssim 0.5\,{\rm pc}$ (von Fellenberg et al., 2022), which implies a mean stellar density of $n_{\star}\sim 382\,{\rm pc^{-3}}$ and a mean stellar distance of $d_{\star}\sim 2.8\times 10^{4}\,{\rm AU}$. In the stellar system analogous to IRS 13E that potentially hosts an IMBH, the mean stellar density within $20\,{\rm mpc}$ is $n_{\star}\sim 1.8\times 10^{5}\,{\rm pc^{-3}}$, which is three orders of magnitude larger than around Sgr A*. The mean distance between stars is $d_{\star}\sim 3.7\times 10^{3}\,{\rm AU}$, which is one order of magnitude smaller than around Sgr A*. This implies a much more frequent and vigorous wind-wind interaction in compact stellar systems around the IMBH in comparison with the Sgr A* stellar environment.
@@ 2.3 Computational grid and boundary conditions
Most simulations are performed on a 3D Cartesian grid with a resolution of $256^{3}$ grid cells, covering a $(0.06\,\text{pc})^{3}$ volume. We employ 5 levels of nested mesh refinement at points $r=r_{0}/2^{n}$, where $r_{0}$ is the box edge length and $n$ indicates the refinement level. This approach effectively doubles the simulation’s resolution with each halving of the radial distance. Therefore, the edge length of the smallest grid cell is $\approx 7\times 10^{-6}$ pc. The global character of the fluid variables is relatively independent of resolution, as we have tested. We ran simulations at $64^{3}$ and $128^{3}$ and found qualitatively similar results. However, we found $256^{3}$ necessary to better resolve the cold clumps that form in some of our simulations. For the simulations with chemical abundance model I, where cold clumps do not form, we use a base resolution of 1283. The inner boundary is equal to twice the edge of the smallest grid cell, $r_{\text{in}}\approx 1.4\times 10^{-5}\,\text{pc}\approx 9765\,r_{\text{g}}$, to ensure proper variable reconstruction for cells bordering the inner boundary. We let the simulations evolve in time up to $t\approx 2$ kyr, or about 1.5 orbital periods for the most distant star, E7.
The code evolves the conservative variables of mass density, momentum density, and total energy density. Consequently, $\rho$ and $P$ can potentially reach unphysical (i.e. negative) values in regions with low mass. To prevent code failure in such an event, we utilize floors on density and pressure such that if $\rho<\rho_{\rm floor}$, we set $\rho=\rho_{\rm floor}$, and similarly if $P<P_{\rm floor}$ we set $P=P_{\rm floor}$. We adopt the values $\rho_{\rm floor}=10^{-7}M_{\odot}\,{\rm pc}^{-3}$ and $P_{\rm floor}=10^{-10}M_{\odot}\,{\rm pc}^{-1}{\rm kyr}^{-2}$. Furthermore, we impose a minimum temperature of $10^{4}$ K, providing an additional, density-dependent floor on pressure. The floors are activated sufficiently rarely and, therefore, do not affect our results.
The region within the inner boundary is set to have a floored density, pressure, and zero velocity. This approach allows the material to fall into the ‘black hole’ while the unphysical boundary affects only a few cells outside of $r_{\text{in}}$. The outer boundary condition is set to outflow in all directions.
@@ 2.4 Model Approximations
Our study is the first of its kind to investigate wind-fed accretion onto an IMBH embedded in a compact stellar cluster using a 3D MHD framework. Given the lack of prior studies in this specific parameter regime, we carefully selected our model parameters to ensure a physically realistic, yet computationally feasible approach. Our goal is to isolate the dominant physical mechanisms governing the accretion process while minimizing additional complexities that could obscure fundamental trends. However, certain approximations are necessary, and these should be considered when interpreting our results.
First, we assume that the stars follow fixed, circular orbits with zero eccentricity. In reality, stellar orbits in a dense cluster are expected to have a range of eccentricities due to dynamical interactions, which could introduce periodic variations in wind densities and collision velocities. Over longer timescales, eccentric orbits may enhance variability in the wind structure and introduce episodic increases in the accretion rate as stars approach the pericenter. Additionally, previous studies have shown that when winds originate from stars on eccentric orbits, the trailing portions of the wind can be captured by the black hole, leading to angular momentum accumulation and the possible formation of a rotationally supported accretion disk (Cuadra et al., 2008). By restricting our setup to circular orbits, we may be underestimating the degree to which wind-fed accretion transitions from a quasi-spherical inflow to a more disk-like structure. Furthermore, the orbits in IRS 13E likely evolve in time. However, given that our simulations cover only a relatively short timescale of $\sim 2$ kyr, the effects of orbital evolution should be minimal.
@@ 3 Results
The differences are also visible in Fig. 2, which shows plane-parallel cuts at $y=0$ of the simulation box (note that there is no axis of symmetry) of electron number density $n_{\mathrm{e}}=\rho/(\mu_{\mathrm{e}}m_{\mathrm{p}})$, accretion rate, $\dot{M}=4\pi r^{2}\rho v_{\mathrm{r}}$, and temperature $T=\mu m_{\mathrm{p}}P/(\rho\mathrm{k_{B}})$, where $m_{\rm p}$ is the proton mass, $v_{\rm r}$ radial velocity, and $\mathrm{k_{B}}$ is the Boltzmann constant, for the two discussed scenarios (left and right columns), as well as Model II with magnetized winds (middle column). In the lower-metallicity case (Fig. 2, left column), reduced radiative cooling efficiency prevents the formation of cool, dense clumps within the wind. Instead, this scenario facilitates the development of high-temperature ($\sim 10^{7}$ K) outflow channels that extend beyond the cluster and originate where the winds collide. In the simulation with higher metallicity of the stellar winds (Fig. 2, right column), the radiative cooling is more efficient, leading to the formation and ejection of cool, dense clumps. There is no sign of an accretion disk forming in any of our simulations, instead, accretion of matter happens primarily through close encounters of the stellar winds and the central black hole. This is true even for simulations with a disk-like distribution of stars (not shown).
Compared to the pure HD case, the Model II simulation with magnetic fields (Fig. 2, middle column) shows minimal differences, as magnetic fields do not appear to significantly alter the global dynamics.

### v2

@@ 2 Methods
We conduct our simulations with athena++, a 3D grid-based scheme that solves the equations of conservative MHD (Stone et al., 2020). Our simulations are based on the HD stellar wind model described in Ressler et al. (2018) and extended in Ressler et al. (2019a, b) to include magnetized winds. The stellar winds are treated as source terms in mass, momentum, energy, and magnetic field moving on fixed Keplerian orbits. The hydrodynamic properties of the stellar winds characterized their mass-loss rates, $\dot{M_{\mathrm{w}}}$, and terminal stellar-wind velocities, $v_{\mathrm{w}}$. Furthermore, we include magnetization of the winds in some of our simulations. The magnetic fields are solely toroidal with respect to the spin axes of the stars (unknown but chosen randomly), and their magnitude is expressed through the $\beta_{\mathrm{w}}$ parameter, defined as the ratio between the wind ram pressure and its magnetic pressure at the equator. In the general scenario involving magnetized winds, the simulation solves the following set of equations:
@@
$\frac{\partial\mathbf{B}}{\partial t}-\nabla\times(\mathbf{v}\times\mathbf{B})$ $=\nabla\times(\mathbf{\tilde{E}}_{\mathrm{w}}),$
where $\rho$ is the mass density, $\mathbf{v}$ is the velocity vector, $P_{\mathrm{tot}}=P+B^{2}/2$ is the total pressure with both thermal and magnetic contributions, $\mathbf{I}$ and $\hat{\mathbf{r}}$ represent the unit matrix and the unit vector respectively, $\mathbf{B}$ is the magnetic field vector, $E=1/2\rho v^{2}+P/(\gamma-1)+B^{2}/2$ represents the total energy with $\gamma=5/3$ being the non–relativistic adiabatic index of the gas, $q_{-}$ is the cooling rate per unit volume due to radiative losses, $f$ is the fraction of the cell by volume contained in the wind, $\dot{\rho}_{\mathrm{w}}=\dot{M}_{\mathrm{w}}/V_{\mathrm{w}}$ with $V_{\mathrm{w}}=4/3(\pi r_{\rm w}^{3})$, $\mathbf{v}_{\mathrm{w,net}}$ is the wind speed in the fixed frame of the grid, $\langle\rangle$ denotes a volume average of a quantity over the cell, $\dot{E}_{\mathrm{B}}$ is the magnetic energy source term generated by the winds, $\mathbf{\tilde{E}}_{\mathrm{w}}$ is the average of the wind source electric field $\mathbf{E}_{\mathrm{w}}$ over the appropriate cell edge. In the simulation domain, each ‘wind’ has a radius $r_{\mathrm{w}}\approx 2\sqrt{3}\Delta x$, where $\Delta x$ is the edge length of the cell which contains the center of the star. Note that here and throughout we use Lorentz-Heaviside units so that a factor of $\sqrt{4\mathrm{\pi}}$ has been absorbed into the definition of $\mathbf{B}$ and the magnetic pressure is $P_{B}=B^{2}/2$.
@@ 2.1 Cooling function and chemical abundance models
Due to the proximity of the stellar winds, they interact and shock-heat, resulting in significant radiative losses due to optically thin bremsstrahlung and line cooling. To calculate the cooling function, $\Lambda$, for arbitrary hydrogen, helium, and metal mass fractions ($X$, $Y$, $Z$ respectively), we first determine the cooling curve for the solar abundances $X_{\odot}=0.7491,\,Y_{\odot}=0.2377,\,Z_{\odot}=0.0133$, presented in Lodders (2003). The calculation is done using the spectral analysis code spex (Kaastra et al., 1996), following the methodology outlined in Schure et al. (2009) to calculate the separate contributions of individual elements to the cooling function,
@@
Here, we assume that oxygen contributes most of the mass in metals and treat both $\mu_{\rm e}$ and $\mu$ as constants. The cooling function $\Lambda$ is implemented by calculating the cooling rate per unit volume using a piece-wise power law approximation over a range of temperatures from $10^{4}$ to $10^{9}$ K. For additional details on the computation and implementation of the cooling function in the simulations, refer to Sect. 2.2 in Ressler et al. (2018).
The real chemical composition of WR stars in compact stellar clusters remains uncertain. We, therefore, use three different chemical abundance models of WR stars used previously in Galactic center simulations (Cuadra et al., 2008; Calderón et al., 2016; Russell et al., 2017). First, we consider hydrogen-depleted stars ($X=0$) with higher metallicity ($Z=3Z_{\odot}$), referred to as Model I in Table 1. This model motivated the fact that WR stars typically lose their outer hydrogen later in previous stages of stellar evolution. Additionally, we explore two alternative chemical abundance models: one with a higher hydrogen content (Model II), which aligns with a realistic expectation that the stars are not completely starved of hydrogen, and another with significantly elevated metallicity (Model III), to investigate effects such as enhanced cooling efficiency and the potential formation of cool clumps due to stellar wind interactions. The accretion flow in Galactic center simulations can be quite sensitive to these mass fractions, as was recently shown by Calderón et al. (2025).
@@ 2.2 Simulation Setup
Our setup is loosely based on the stellar association IRS 13E within the Milky Way nuclear stellar cluster near Sgr A*. The six WR stars in our simulation are placed at fixed distances from the central black hole, matching the observed projected distances of these stars from the hypothetical IMBH with an approximate mass of $\sim 3\times 10^{4}M_{\odot}$, associated with the IRS 13E3 source (see Peißker et al., 2024, for more details). That said, we use IRS 13E primarily as a reference for the stellar properties, taking a more flexible approach to other key properties, such as the spatial distribution of the stars, their chemical abundance models, and the inclusion of wind magnetization.
@@
Sim. Chem. model Stellar distr. $\beta_{\mathrm{w}}$ Res. SMR
1 I sph. A $\infty$ $128^{3}$ 5
2 II sph. B $\infty$ $256^{3}$ 5
3 II sph. B 100 $256^{3}$ 5
4 III sph. A $\infty$ $256^{3}$ 5
5 I disk $\infty$ $128^{3}$ 5
6 III disk $\infty$ $256^{3}$ 5
  • Notes. A and B in the spherical stellar distributions denote different random realizations.
Figure 1: 3D volume rendering of the density of the simulated stellar cluster. Left: Colliding stellar winds 620 years into the Model II spherical simulation. Right: Outbursts of cool clumps 815 years into the Model III spherical simulation, driven by the gas’s higher metal content, which enhances cooling efficiency and amplifies the development of the thin-shell instability.
Table 3: Orbital parameters of the simulated stars.
Source ID $r$ $10^{-3}$pc] $v$ [km s-1] $P$ [yr]
@@ 6 III disk $\infty$ $256^{3}$ 5
  • Notes. Setup inspired by the IRS 13E association; $r$ from projected separations to the putative IMBH (IRS 13E3). Source IDs from Peißker et al. (2023).
Table 2 provides a summary of the simulation runs analyzed in this study. In our simulation, we assume that all the stars are identical WR stars, each characterized by a wind mass-loss rate of $\dot{M}_{\text{w}}=10^{-5}\,M_{\odot}\,\mathrm{yr}^{-1}$ and a stellar wind speed of $v_{\mathrm{w}}\penalty 10000\ =\penalty 10000\ 1000\,\mathrm{km\,s}^{-1}$, which are typical for the WR stars in the Galactic center region (Krabbe et al., 1995; Moultaka et al., 2005; Wang et al., 2020). These winds are moving on fixed Keplerian orbits around the center of the simulation domain. The stars are on circular orbits, with the orbital radius taken from projected distances (as listed in Table 3) and initial positions along the orbits generated randomly. We randomize the positions for a spherically isotropic stellar distribution entirely, except for the fixed orbital distance. In contrast, for a disk-like distribution, the inclination angles of the stellar orbits are constrained to $i=\pm 15^{\circ}$.
The setup summarized in Table 3 indicates that for the mean stellar distance of $r\sim 10\,{\rm mpc}$, the orbital velocity is $v=(GM_{\bullet}/r)^{1/2}\simeq 113.6(M_{\bullet}/3\times 10^{4}\,M_{\odot})^{1/2}(r/10\,{\rm mpc})^{-1/2}\,{\rm km\,s^{-1}}$, implying a stellar-wind velocity to orbital velocity ratio of $v_{\rm w}/v\sim 8.8$. For the stars bound to Sgr A*, the orbital velocities within the clockwise disc are of the order of $v\sim 415\,(M_{\bullet}/4\times 10^{6}\,M_{\odot})^{1/2}(r/0.1\,{\rm pc})^{-1/2}\,{\rm km\,s^{-1}}$ and the ratio is $v_{\rm w}/v\sim 2.4$, which is $\sim 4$ times smaller than for the case of adopted orbits around the IMBH. The difference in $v_{\rm w}/v$ clearly motivates the exploration of the wind-fed accretion regime in compact stellar systems around the IMBH.
Another motivation is the much higher mean stellar density than young massive stars around Sgr A*. Around Sgr A*, there are about $\sim 200$ young massive stars within $r\lesssim 0.5\,{\rm pc}$ (von Fellenberg et al., 2022), which implies a mean stellar density of $n_{\star}\sim 382\,{\rm pc^{-3}}$ and a mean stellar distance of $d_{\star}\sim 2.8\times 10^{4}\,{\rm AU}$. In the stellar system analogous to IRS 13E that potentially hosts an IMBH, the mean stellar density within $20\,{\rm mpc}$ is $n_{\star}\sim 1.8\times 10^{5}\,{\rm pc^{-3}}$, which is three orders of magnitude larger than around Sgr A*. The mean distance between stars is $d_{\star}\sim 3.7\times 10^{3}\,{\rm AU}$, one order of magnitude smaller than around Sgr A*. This implies a much more frequent and vigorous wind-wind interaction in compact stellar systems around the IMBH compared to the Sgr A* stellar environment.
@@ 2.3 Computational grid and boundary conditions
Most simulations are performed on a 3D Cartesian grid with a $256^{3}$ grid cell resolution, covering a $(0.06\,\text{pc})^{3}$ volume. We employ five levels of nested mesh refinement at points $r=r_{0}/2^{n}$, where $r_{0}$ is the box edge length and $n$ indicates the refinement level. This approach effectively doubles the simulation’s resolution with each halving of the radial distance. Therefore, the edge length of the smallest grid cell is $\approx 7\times 10^{-6}$ pc. The global character of the fluid variables is relatively independent of resolution, as we have tested. We ran simulations at $64^{3}$ and $128^{3}$ and found qualitatively similar results. However, we found $256^{3}$ necessary to resolve the cold clumps in some simulations. For the simulations with the chemical abundance model I, where cold clumps do not form, we use a base resolution of 1283. The inner boundary is equal to twice the edge of the smallest grid cell, $r_{\text{in}}\approx 1.4\times 10^{-5}\,\text{pc}\approx 9765\,r_{\text{g}}$, to ensure proper variable reconstruction for cells bordering the inner boundary. We let the simulations evolve in time up to $t\approx 2$ kyr, or about 1.5 orbital periods for the most distant star, E7.
The code evolves the conservative variables of mass density, momentum density, and total energy density. Consequently, $\rho$ and $P$ can reach unphysical (i.e., negative) values in regions with low mass. To prevent code failure in such an event, we utilize floors on density and pressure such that if $\rho<\rho_{\rm floor}$, we set $\rho=\rho_{\rm floor}$, and similarly if $P<P_{\rm floor}$ we set $P=P_{\rm floor}$. We adopt the values $\rho_{\rm floor}=10^{-7}M_{\odot}\,{\rm pc}^{-3}$ and $P_{\rm floor}=10^{-10}M_{\odot}\,{\rm pc}^{-1}{\rm kyr}^{-2}$. Furthermore, we impose a minimum temperature of $10^{4}$ K, providing an additional, density-dependent floor on pressure. The floors are activated sufficiently rarely and, therefore, do not affect our results.
We set the region within the inner boundary to have a floored density, pressure, and zero velocity. This approach allows the material to fall into the ‘black hole’ while the unphysical boundary affects only a few cells outside of $r_{\text{in}}$. The outer boundary condition is set to outflow in all directions.
@@ 2.4 Model Approximations
Our study is the first to investigate wind-fed accretion onto an IMBH embedded in a compact stellar cluster using a 3D MHD framework. Given the lack of prior studies in this specific parameter regime, we carefully selected our model parameters to ensure a physically realistic, yet computationally feasible approach. We aim to isolate the dominant physical mechanisms governing the accretion process while minimizing additional complexities that could obscure fundamental trends. However, certain approximations are necessary to consider when interpreting our results.
First, we assume the stars follow fixed, circular orbits with zero eccentricity. In reality, stellar orbits in a dense cluster are expected to have a range of eccentricities due to dynamical interactions, which could introduce periodic variations in wind densities and collision velocities. Over longer timescales, eccentric orbits may enhance variability in the wind structure and introduce episodic increases in the accretion rate as stars approach the pericenter. Additionally, previous studies have shown that when winds originate from stars on eccentric orbits, the trailing portions of the wind can be captured by the black hole, leading to angular momentum accumulation and the possible formation of a rotationally supported accretion disk (Cuadra et al., 2008). By restricting our setup to circular orbits, we may underestimate how wind-fed accretion transitions from a quasi-spherical inflow to a more disk-like structure. Furthermore, the orbits in IRS 13E are likely to evolve over time. However, given that our simulations cover only a relatively short timescale of $\sim 2$ kyr, the effects of orbital evolution should be minimal.
@@ 3.1 Flow morphology
The differences are also visible in Fig. 2, which shows plane-parallel cuts at $y=0$ of the simulation box (note that there is no axis of symmetry) of electron number density $n_{\mathrm{e}}=\rho/(\mu_{\mathrm{e}}m_{\mathrm{p}})$, accretion rate, $\dot{M}=4\pi r^{2}\rho v_{\mathrm{r}}$, and temperature $T=\mu m_{\mathrm{p}}P/(\rho\mathrm{k_{B}})$, where $m_{\rm p}$ is the proton mass, $v_{\rm r}$ radial velocity, and $\mathrm{k_{B}}$ is the Boltzmann constant, for the two discussed scenarios (left and right columns), as well as Model II with magnetized winds (middle column). In the lower-metallicity case (Fig. 2, left column), reduced radiative cooling efficiency prevents the formation of cool, dense clumps within the wind. Instead, this scenario facilitates the development of high-temperature ($\sim 10^{7}$ K) outflow channels that extend beyond the cluster and originate where the winds collide.
In the simulation with higher metallicity of the stellar winds (Fig. 2, right column), the radiative cooling is more efficient, leading to the formation and ejection of cool, dense clumps. There is no sign of an accretion disk forming in our simulations; instead, accretion of matter happens primarily through close encounters of the stellar winds and the central black hole. This is true even for simulations with a disk-like distribution of stars (not shown).
Compared to the pure HD case, the Model II simulation with magnetic fields (Fig. 2, middle column) shows minimal differences, as magnetic fields do not appear to alter the global dynamics significantly.